# DNOI-4DRO: Deep 4D Radar Odometry with Differentiable Neural-Optimization Iterations

### v1

@@ Abstract
1 Introduction
Odometry is critical to autonomous driving systems, particularly in GPS-denied environments. It involves using consecutive images or point clouds to estimate the relative pose transformation between two frames. Most existing odometry methods focus primarily on 2D visual odometry droid ; deeppatch ; multi ; mambavo ; leap or 3D LiDAR pwclo ; lonet ; efficientlo ; translo ; delo ; nerfloam ; dslo . However, the inherent characteristics of cameras and LiDARs make these methods vulnerable to challenging weather conditions such as rain, snow, or fog, compromising their robustness in these scenarios.
4D millimeter-wave radar, as an emerging automotive sensor, presents distinct advantages, including robustness under challenging weather and illumination conditions, the ability to measure object velocities and cost-effectiveness ratrack ; milliflow ; radarocc ; rcbevdet ; transloc4d ; talk2radar . These advantages have driven extensive research into 4D Radar Odometry (4DRO). Previous 4D radar odometry works can be classified into two categories: classical methods 4dradarslam ; li20234d ; efear and learning-based methods 4drvonet ; 4dronet . Classical 4DRO systems typically consist of point cloud association and nonlinear optimization. Nonlinear optimization is critical in enhancing accuracy by integrating continuous sensor observations to optimize long-range trajectories. However, these methods have limited robustness because of the extracted features’ poor quality and low resolution 4dradarslam ; li20234d . With the development of deep learning, some methods 4dronet ; cao attempt to utilize learning-based methods for 4D radar pose estimation. These methods extract features from two point clouds using a feature extraction network, perform point cloud matching through a feature association network, and finally estimate the pose using a decoding network. Although this direct pose regression method, which does not rely on explicit optimization, can sometimes be more robust, its accuracy often shows limitations across various environments.
@@ 1 Introduction
(1) We propose DNOI-4DRO, a novel framework for 4D radar odometry. By exploiting the relationship between point motion and pose, DNOI-4DRO couples a geometric optimizer with a neural network to construct a fully end-to-end architecture. (2) We specially design an efficient radar feature extractor comprising a dual-stream 4D radar backbone, which extracts radar features with two representations and is further enhanced by a global Transformer. (3) Finally, our method is demonstrated on VoD vod and Snail-Radar snailradar datasets. Our method outperforms all classical and learning-based 4D radar odometry methods, achieving competitive results with A-LOAM, which uses LiDAR point clouds as input.
2 Related Work
2.1 Classical 4D Radar Odometry
@@
Due to the similarity in spatial representation between 4D radar and LiDAR point clouds, most existing classical 4D radar odometry methods either directly adopt LiDAR odometry methods or modify them accordingly. Iterative Closest Point (ICP) is the most widely used LiDAR odometry method, computing relative pose transformations by minimizing the Euclidean distance between two point clouds. Based on error measurement, ICP can be categorized into ICP-point2point p2picp and ICP-point2plane p2plicp , aiming to shorten point-to-point and point-to-plane distances. Generalized ICP (GICP) gicp further intends to combine the advantages of both. Building on GICP, 4DRadarSLAM 4dradarslam models the spatial probabilistic distribution of radar points based on uncertainties in range and azimuth angle measurements. It proposes an Adaptive Probability Distribution-GICP (APDGICP) method based on the modeling results. RIV-SLAM rivslam further refines this by considering the impact of the angle of arrival on radar measurement uncertainty to achieve a more accurate spatial probabilistic distribution of radar points. In addition to ICP, the Normal Distribution Transform (NDT) ndt is a widely used LiDAR odometry method. NDT represents point clouds with Gaussian distributions, converting pose estimation into an optimization problem that minimizes the Gaussian distribution error between the source and target point clouds. Expanding upon NDT, Li et al. li20234d integrate the measurement uncertainties of 4D radar points into the calculations for the mean and variance of the normal distribution, aiming to reduce degradation effects caused by sparse 4D radar point clouds. Additionally, 4D iRIOM 4driom introduces a point cloud registration method based on a distribution-to-multi-distribution distance metric, effectively addressing the sparsity issue of 4D radar point clouds and enhancing pose estimation robustness.
@@
Our method draws on the cost function modeling approach of the ICP-point2point p2picp , a simpler optimization problem that avoids extracting complex representations like line and plane features. Instead of relying on nearest-neighbor queries to find matching points, our method uses a neural network to estimate the point motion field. This approach mitigates the challenging non-bijective correspondence between two point clouds caused by the sparsity and noise inherent in 4D radar point clouds.
2.2 Deep 4D Radar Odometry
@@
Deep learning has made significant strides in visual and LiDAR odometry, yet 4D radar odometry remains a challenging problem. 4DRONet 4dronet proposes a sliding-window-based hierarchical optimization method to estimate and refine poses in a coarse-to-fine manner. SelfRO 10422466 introduces a self-supervised 4D radar odometry method that uses radar point velocity information to construct a novel velocity-aware loss function, effectively guiding network training. CMFlow cmflow introduces a cross-modal supervision approach for estimating 4D radar scene flow, simultaneously providing pose estimation as an intermediate output. CAO-RONet cao introduces a sliding-window-based optimizer that leverages historical priors to enable coupled state estimation and correct inter-frame matching errors. Although these methods show promising results, they often suffer from limited accuracy across diverse environments, and their interpretability remains constrained.
@@
Moreover, we are inspired by the “Differentiable Recurrent Optimization-Inspired Design” proposed by DROID-SLAM droid for visual odometry. The work combines iterative visual correspondence updates with differentiable bundle adjustments to optimize pose estimation. This concept has also been applied to optimizing pixel-level 3D motion 3draft . However, both works are based on RAFT raft , which performs optical flow updates and optimizes variables in 2D space. This paper applies the idea to 3D space and a neural-optimization iteration module specifically for 3D point cloud odometry is being built for the first time.
3 DNOI-4DRO
@@
We propose a backbone network for pose estimation between two 4D radar point clouds and then construct a 4DRO system upon it. We provide an overview in Fig. 1. The backbone network firstly extracts robust point and context features for each 4D radar point cloud, as introduced in Sec. 3.1. Then, the feature correlation volume is constructed using point features, which will be described in Sec. 3.2. Next, the differentiable neural-optimization iteration operator is introduced in Sec. 3.3 for pose estimation and refinement. Finally, the network outputs the pose $\mathbf{T}^{1}\in\mathbb{R}^{4\times 4}$ and $\mathbf{T}^{2}\in\mathbb{R}^{4\times 4}$ of the two 4D radar point clouds in the world coordinate system.
@@
Figure 1: Overview of our backbone. (1) The feature and context extractors encode the point and context features of the input point cloud, respectively. (2) The feature correlation module constructs an all-pair correlation volume by calculating the matrix dot product of two-point features. (3) In each iteration, the differentiable neural-optimization iteration operator uses the pose estimated in the previous iteration to look up correlation features from the correlation volume, which are then processed through a GRU gru to generate a point motion field. The point motion field is fed into a least-squares-based optimization layer, where the pose is updated based on geometric constraints. After multiple iterations, the network outputs the predicted pose.
3.1 Feature Extraction
@@
We employ two radar feature extraction networks with identical architectures but distinct functions to extract the point and context features from the 4D radar point cloud. The context features are provided as input to the recurrent iteration operator, whereas the point features are used to evaluate the point similarity using a dot product. To obtain fine-grained 4D radar point features, we propose a dual-stream 4D radar encoding network that integrates multi-scale geometric feature extraction, clustering-based class-aware feature extraction, and a global Transformer module, as illustrated in Fig. 2. This network offers a more comprehensive feature representation for each 4D radar point, thereby significantly improving the robustness of odometry across diverse environments.
3.1.1 Multi-scale Geometric Feature Extraction
@@
Given the 4D radar point clouds $\mathbf{P}=\{\mathbf{p}_{i}|\mathbf{p}_{i}\in{\mathbb{R}^{3}}\}_{i=1}^{N}$, we employ four parallel Set Abstraction (SA) layers pointnet++ to encode the multi-scale geometric features. Each SA layer uses a different grouping radius to address the issue of non-uniform density in radar point clouds. At each scale $s$, each radar point $\mathbf{p}_{i}$ generates a local geometric feature $\mathbf{f}_{i,s}^{\text{geo}}$. The geometric features across all scales are concatenated to form the multi-scale geometric feature representation of the 4D radar point, $\mathbf{f}_{i}^{\text{geo}}=\text{concat}(\{\mathbf{f}_{i,s}^{\text{geo}}\}_{s=0}^{3})$.
3.1.2 Clustering-based Class-aware Feature Extraction
Inspired by Context Clusters ma2023image , which proposes a generic clustering-based visual backbone viewing images as a set of points, we extend it and propose a novel clustering-based class-aware 4D radar point feature extraction method, consisting primarily of three components: 4D radar point clustering, class-aware feature aggregation, and feature dispatching.
4D Radar Point Clustering: We first use PointNet pointnet to map the 4D radar points to a high-dimensional feature space $\mathbf{f}_{i}^{\text{p}}$ for similarity computation. Next, we apply Farthest Point Sampling (FPS) pointnet++ to evenly select $C$ center points in space, and the center feature $\mathbf{f}^{\text{ct}}$ is computed by averaging its $K$ nearest points. We then calculate the pairwise cosine similarity matrix $\mathbf{S}\in\mathbb{R}^{C\times N}$ between each point and the set of center points. After that, we allocate each point to the most similar center, resulting in $C$ clusters.
Class-aware Feature Aggregation: Following ma2023image , we dynamically aggregate all radar point features within the same cluster based on the similarities to the cluster center. Given the cluster which contains $M$ radar points around the $j$-th cluster center, the aggregated feature $\mathbf{f}_{j}^{\text{a}}$ is calculated by:
$\mathbf{f}_{j}^{\text{a}}=\frac{1}{\mathbf{A}}\left(\mathbf{f}_{j}^{\text{ct}}+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{m}^{\text{p}}\right),\quad\mathbf{A}=1+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta),$ (1)
@@ 3.1.2 Clustering-based Class-aware Feature Extraction
3.1.3 Global Transformer
The global transformer employs an attention mechanism attention across the entire radar point cloud to capture the long-range correlation of points. For each point $\mathbf{p}_{i}\in\mathbf{P}$, the global transformer applies an attention mechanism to all other points $\mathbf{p}_{j}\in\mathbf{P}$:
@@ 3.1.3 Global Transformer
[…]
3.2 Feature Correlation
@@
3.3 Differentiable Neural-Optimization Iteration Operator
The core of our approach is the differentiable neural-optimization iteration operator, which contains three key parts, as illustrated in Fig. 3: (1) an adaptive patch-to-patch correlation lookup method that retrieves correlation features from the precomputed correlation volume; (2) a GRU gru that predicts updates to the point motion flow along with associated confidence weights; and (3) an All-point Motion-only Bundle Adjustment layer (AMBA) computes pose that aligns with the newly predicted point motion flow and confidence.
@@ 3.3 Differentiable Neural-Optimization Iteration Operator
3.3.1 Adaptive Patch-to-Patch Correlation Look-up Layer
@@
$\mathbf{P}^{12}_{i}$ represents the coordinates of points $\mathbf{P}^{1}$ mapped into $\mathbf{P}^{2}$ using the estimated pose. Then, we incorporate the Patch-to-Patch correlation concept from HALFlow wang2021hierarchical into the point correlation look-up mechanism of PV-RAFT pvraft to learn fine-grained correlation features of sparse 4D radar points, thereby guiding the right flowing direction of point motion. Specifically, for each point $\mathbf{p}_{i,j}^{12}$ in $\mathbf{P}^{12}_{i}$, we first retrieve the initial correlation features from the correlation volume using nearest neighbor queries. Subsequently, based on the 3D local geometric structure of $\mathbf{p}_{i,j}^{12}$ and its neighboring points in $\mathbf{P}^{12}_{i}$, we adaptively weight the initial correlation features of these neighbors to obtain the final correlation feature $\mathbf{cf}_{i,j}^{12}$.
3.3.2 Point Motion Flow Update and Confidence Weight Prediction
@@
The motion features $\mathbf{MF}^{12}_{i}$ is used as the input to the GRU gru to update the current hidden state $\mathbf{h}^{12}_{i}$. Instead of predicting updates to the pose directly, we predict updates in the space of point motion flow fields. We map the updated hidden state through two additional MLP layers to produce two outputs: (1) a revision flow field $\Delta\mathbf{FL}^{12}_{i}\in\mathbb{R}^{N\times 3}$ and (2) associated confidence $\mathbf{w}^{12}_{i}\in\mathbb{R}_{+}^{N\times 3}$. The revision $\Delta\mathbf{FL}^{12}_{i}$ is a correction term predicted by the network to correct errors in the all-point correspondence field. We denote the corrected correspondence as $\mathbf{P}^{12*}_{i}=\mathbf{P}^{12}_{i}+\Delta\mathbf{FL}^{12}_{i}$. The confidence term is further processed with a sigmoid operation to constrain its values within the $(0,1)$. This term mitigates the adverse impact of inaccurate point motion flow estimation caused by occlusion, moving objects and noisy points on pose estimation accuracy.
3.3.3 All-point Motion-only Bundle Adjustment Layer
@@
$E(\mathbf{T}^{1},\mathbf{T}^{2})=\left\|r(\mathbf{T}^{1},\mathbf{T}^{2})\right\|^{2}_{\Sigma_{12}}=\left\|\mathbf{P}^{12*}_{i}-\mathbf{T}^{2}\cdot(\mathbf{T}^{1})^{-1}\cdot\mathbf{P}^{1}\right\|^{2}_{\Sigma_{12}},$ (7)
where $\left\|\cdot\right\|^{2}_{\Sigma}$ is the Mahalanobis distance which weights the error terms based on the confidence weights $\mathbf{w}^{12}_{i}$. In this paper, we optimize the pose using the Gauss-Newton algorithm, based on the special Euclidean Lie algebra $\text{se}(3)=\left\{\xi=\left(\begin{array}[]{c}\rho\\ \phi\end{array}\right),\ \rho\in\mathbb{R}^{3},\ \rho\in\text{so}(3)\right\}$. First, we compute the Jacobian matrices $\mathbf{J}^{1}\in\mathbb{R}^{N\times 3\times 6}$ and $\mathbf{J}^{2}\in\mathbb{R}^{N\times 3\times 6}$ of the residual term $r(\mathbf{T}^{1},\mathbf{T}^{2})$ with respect to $\mathbf{T}^{1}$ and $\mathbf{T}^{2}$ using the perturbation model. For $\mathbf{T}^{2}$, we apply the perturbation $\text{Exp}(\xi^{2})$ on the left, which results in:
$\mathbf{J}^{2}=\frac{\partial r}{\partial\xi^{2}}=-\lim_{\xi^{2}\to 0}\frac{\text{Exp}(\xi^{2})\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{2}}=-\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right],$ (8)
where $\text{Exp}(\cdot)$ represents the exponential map from Lie algebra to Lie group, and $\left(\cdot\right)^{\wedge}$ denotes the antisymmetric matrix operation. For $\mathbf{T}^{1}$, we similarly apply the perturbation $\text{Exp}(\xi^{1})$ and use the adjoint operator and chain rule of differentiation, which results in:
$\mathbf{J}^{1}=\frac{\partial r}{\partial\xi^{1}}$ $=-\lim_{\xi^{1}\to 0}\frac{\text{Exp}\left(-\text{Adj}_{\mathbf{T}^{12}}\xi^{1}\right)\circ\mathbf{T}^{12}\circ\mathbf{P}^{1}-\mathbf{T}^{12}\circ\mathbf{P}^{1}}{\xi^{1}}$ (9)
$=\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right]\circ\text{Adj}_{\mathbf{T}^{12}},$
where $\text{Adj}_{\left(\cdot\right)}$ denotes the adjoint operator. We concatenate the Jacobian matrices $\mathbf{J}^{1}$ and $\mathbf{J}^{2}$ to obtain the Jacobian matrix $\mathbf{J}\in\mathbb{R}^{N\times 3\times 12}$ of the residual term concerning the optimization variables.
Then, we linearizes the residual terms, and solves for the update:
$\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J}\Delta\xi=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}),$ (10)
where $\Delta\xi=(\Delta\xi^{1},\Delta\xi^{2})$ represents the pose update in the Lie algebra. Eqn.10 can be rewritten as a linear system:
$\mathbf{H}\Delta\xi=\mathbf{b},\quad\mathbf{H}=\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J},\quad\mathbf{b}=-\mathbf{J}^{T}r(\mathbf{T}^{1},\mathbf{T}^{2}).$ (11)
@@
4 Training and Implementation
DNOI-4DRO is implemented in PyTorch, with the LieTorch extension teed2021tangent used to perform backpropagation in the tangent space of all group elements.
4.1 Constructing Frame Graph

### v2

@@ Abstract
Introduction
Odometry is critical to autonomous driving systems, particularly in GPS-denied environments. It involves using consecutive images or point clouds to estimate the relative pose transformation between two frames. Most existing odometry methods focus primarily on 2D visual odometry (droid; deeppatch; multi; mambavo; leap) or 3D LiDAR (pwclo; lonet; efficientlo; translo; delo; nerfloam; dslo). However, the inherent characteristics of cameras and LiDARs make these methods vulnerable to challenging weather conditions such as rain, snow, or fog, compromising their robustness in these scenarios.
4D millimeter-wave radar, as an emerging automotive sensor, presents distinct advantages, including robustness under challenging weather and illumination conditions, the ability to measure object velocities and cost-effectiveness (ratrack; milliflow; radarocc; rcbevdet; transloc4d; talk2radar). These advantages have driven extensive research into 4D Radar Odometry (4DRO). Previous 4D radar odometry works can be classified into two categories: classical methods (4dradarslam; li20234d; efear) and learning-based methods (4drvonet; 4dronet). Classical 4DRO systems typically consist of point cloud association and nonlinear optimization. Nonlinear optimization is critical in enhancing accuracy by integrating continuous sensor observations to optimize long-range trajectories. However, these methods have limited robustness because of the extracted features’ poor quality and low resolution (4dradarslam; li20234d). With the development of deep learning, some methods (4dronet; cao) attempt to utilize learning-based methods for 4D radar pose estimation. These methods extract features from two point clouds using a feature extraction network, perform point cloud matching through a feature association network, and finally estimate the pose using a decoding network. Although this direct pose regression method, which does not rely on explicit optimization, can sometimes be more robust, its accuracy often shows limitations across various environments.
@@ Introduction
(1) We propose DNOI-4DRO, a novel framework for 4D radar odometry. By exploiting the relationship between point motion and pose, DNOI-4DRO couples a geometric optimizer with a neural network to construct a fully end-to-end architecture. (2) We specially design an efficient radar feature extractor comprising a dual-stream 4D radar backbone, which extracts radar features with two representations. (3) Finally, our method is demonstrated on VoD (vod) and Snail-Radar (snailradar) datasets. Our method outperforms all classical and learning-based 4D radar odometry methods, achieving competitive results with A-LOAM, which uses LiDAR point clouds as input.
Related Work
Classical 4D Radar Odometry
@@
Due to the similarity in spatial representation between 4D radar and LiDAR point clouds, most existing classical 4D radar odometry methods either directly adopt LiDAR odometry methods or modify them accordingly. Iterative Closest Point (ICP) is the most widely used LiDAR odometry method, computing relative pose transformations by minimizing the Euclidean distance between two point clouds. Based on error measurement, ICP can be categorized into ICP-point2point (p2picp) and ICP-point2plane (p2plicp), aiming to shorten point-to-point and point-to-plane distances. Generalized ICP (GICP) (gicp) further intends to combine the advantages of both. Building on GICP, 4DRadarSLAM (4dradarslam) models the spatial probabilistic distribution of radar points based on uncertainties in range and azimuth angle measurements. It proposes an Adaptive Probability Distribution-GICP (APDGICP) method based on the modeling results. RIV-SLAM (rivslam) further refines this by considering the impact of the angle of arrival on radar measurement uncertainty to achieve a more accurate spatial probabilistic distribution of radar points. In addition to ICP, the Normal Distribution Transform (NDT) (ndt) is a widely used LiDAR odometry method. NDT represents point clouds with Gaussian distributions, converting pose estimation into an optimization problem that minimizes the Gaussian distribution error between the source and target point clouds. Expanding upon NDT, Li et al. (li20234d) integrate the measurement uncertainties of 4D radar points into the calculations for the mean and variance of the normal distribution, aiming to reduce degradation effects caused by sparse 4D radar point clouds. Additionally, 4D iRIOM (4driom) introduces a point cloud registration method based on a distribution-to-multi-distribution distance metric, effectively addressing the sparsity issue of 4D radar point clouds and enhancing pose estimation robustness.
@@
Our method draws on the cost function modeling approach of the ICP-point2point (p2picp), a simpler optimization problem that avoids extracting complex representations like line and plane features. Instead of relying on nearest-neighbor queries to find matching points, our method uses a neural network to estimate the point motion field. This approach mitigates the challenging non-bijective correspondence between two point clouds caused by the sparsity and noise inherent in 4D radar point clouds.
Deep 4D Radar Odometry
@@
Deep learning has made significant strides in visual and LiDAR odometry, yet 4D radar odometry remains a challenging problem. 4DRONet (4dronet) proposes a sliding-window-based hierarchical optimization method to estimate and refine poses in a coarse-to-fine manner. SelfRO (10422466) introduces a self-supervised 4D radar odometry method that uses radar point velocity information to construct a novel velocity-aware loss function, effectively guiding network training. CMFlow (cmflow) introduces a cross-modal supervision approach for estimating 4D radar scene flow, simultaneously providing pose estimation as an intermediate output. CAO-RONet (cao) introduces a sliding-window-based optimizer that leverages historical priors to enable coupled state estimation and correct inter-frame matching errors. Although these methods show promising results, they often suffer from limited accuracy across diverse environments, and their interpretability remains constrained.
@@
Moreover, we are inspired by the “Differentiable Recurrent Optimization-Inspired Design” proposed by DROID-SLAM (droid) for visual odometry. The work combines iterative visual correspondence updates with differentiable bundle adjustments to optimize pose estimation. This concept has also been applied to optimizing pixel-level 3D motion (3draft). However, both works are based on RAFT (raft), which performs optical flow updates and optimizes variables in 2D space. This paper introduces the first neural-optimization module for 4D radar-based 3D point cloud odometry by tightly coupling RAFT/DROID-style differentiable iterations with Bundle Adjustment.
DNOI-4DRO
@@
We propose a backbone network for pose estimation between two 4D radar point clouds and then construct a 4DRO system upon it. We provide an overview in Figure 1. The backbone network firstly extracts robust point and context features for each 4D radar point cloud. Then, the feature correlation volume is constructed using point features. Next, the differentiable neural-optimization iteration operator is introduced for pose estimation and refinement. Finally, the network outputs the pose $\mathbf{T}^{1}\in\mathbb{R}^{4\times 4}$ and $\mathbf{T}^{2}\in\mathbb{R}^{4\times 4}$ of the two 4D radar point clouds in the world coordinate system.
@@
Figure 1: Overview of our backbone. (1) The feature and context extractors encode the point and context features of the input point cloud, respectively. (2) The feature correlation module constructs an all-pair correlation volume by calculating the matrix dot product of two-point features. (3) In each iteration, the differentiable neural-optimization iteration operator uses the pose estimated in the previous iteration to look up correlation features from the correlation volume, which are then processed through a GRU (gru) to generate a point motion field. The point motion field is fed into a least-squares-based optimization layer, where the pose is updated based on geometric constraints. After multiple iterations, the network outputs the predicted pose.
Feature Extraction
@@
We employ two radar feature extraction networks with identical architectures but distinct functions to extract the point and context features from the 4D radar point cloud. The context features are provided as input to the recurrent iteration operator, whereas the point features are used to evaluate the point similarity using a dot product. To obtain fine-grained 4D radar point features, we propose a dual-stream 4D radar encoding network that integrates multi-scale geometric feature extraction, clustering-based class-aware feature extraction, and a global Transformer module, as illustrated in Figure 2. This network offers a more comprehensive feature representation for each 4D radar point, thereby significantly improving the robustness of odometry across diverse environments.
Multi-scale Geometric Feature Extraction
@@
Given the 4D radar point clouds $\mathbf{P}=\{\mathbf{p}_{i}|\mathbf{p}_{i}\in{\mathbb{R}^{3}}\}_{i=1}^{N}$, we employ four parallel Set Abstraction (SA) layers (pointnet++) to encode the multi-scale geometric features. Each SA layer uses a different grouping radius to address the issue of non-uniform density in radar point clouds. At each scale $s$, each radar point $\mathbf{p}_{i}$ generates a local geometric feature $\mathbf{f}_{i,s}^{\text{geo}}$. The geometric features across all scales are concatenated to form the multi-scale geometric feature representation of the 4D radar point, $\mathbf{f}_{i}^{\text{geo}}=\text{concat}(\{\mathbf{f}_{i,s}^{\text{geo}}\}_{s=0}^{3})$.
Clustering-based Class-aware Feature Extraction
To address the inherently non-uniform spatial distribution of 4D radar point clouds, we propose a novel clustering-based class-aware 4D radar point feature extraction method, consisting primarily of three components: 4D radar point clustering, class-aware feature aggregation, and feature dispatching.
4D Radar Point Clustering: We first use PointNet (pointnet) to map the 4D radar points to a high-dimensional feature space $\mathbf{f}_{i}^{\text{p}}$ for similarity computation. Next, we apply Farthest Point Sampling (FPS) (pointnet++) to evenly select $C$ center points in space, and the center feature $\mathbf{f}^{\text{ct}}$ is computed by averaging its $K$ nearest points. We then calculate the pairwise cosine similarity matrix $\mathbf{S}\in\mathbb{R}^{C\times N}$ between each point and the set of center points. After that, we allocate each point to the most similar center, resulting in $C$ clusters.
Class-aware Feature Aggregation: Following (ma2023image), we dynamically aggregate all radar point features within the same cluster based on the similarities to the cluster center. Given the cluster which contains $M$ radar points around the $j$-th cluster center, the aggregated feature $\mathbf{f}_{j}^{\text{a}}$ is calculated by:
$\mathbf{f}_{j}^{\text{a}}$ $=\frac{1}{\mathbf{A}}\left(\mathbf{f}_{j}^{\text{ct}}+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta)\cdot\mathbf{f}_{m}^{\text{p}}\right),$ (1)
$\mathbf{A}$ $=1+\sum_{m=1}^{M}\text{sigmoid}(\alpha\mathbf{s}_{jm}+\beta),$ (2)
@@
Global Transformer
The global transformer employs an attention mechanism (attention) across the entire radar point cloud to capture the long-range correlation of points. For each point $\mathbf{p}_{i}\in\mathbf{P}$, the global transformer applies an attention mechanism to all other points $\mathbf{p}_{j}\in\mathbf{P}$:
@@
where $\mathcal{X}^{\text{l}}$ represents the geometric-class joint feature set, $\alpha^{g}(\cdot)$, $\beta^{g}(\cdot)$, and $\gamma^{g}(\cdot)$ are shared learnable linear transformations. $\langle\cdot\rangle$ denotes a scalar product. Linear layer, layer norm, and skip connection are further applied to complete the global transformer module.
Feature Correlation
@@
Differentiable Neural-Optimization Iteration Operator
The core of our approach is the differentiable neural-optimization iteration operator, which contains three key parts, as illustrated in Figure. 3: (1) an adaptive patch-to-patch correlation lookup method that retrieves correlation features from the precomputed correlation volume; (2) a GRU (gru) that predicts updates to the point motion flow along with associated confidence weights; and (3) an All-point Motion-only Bundle Adjustment layer (AMBA) computes pose that aligns with the newly predicted point motion flow and confidence.
@@
Adaptive Patch-to-Patch Correlation Look-up Layer
@@
$\mathbf{P}^{12}_{i}$ represents the coordinates of points $\mathbf{P}^{1}$ mapped into $\mathbf{P}^{2}$ using the estimated pose. Then, we incorporate the Patch-to-Patch correlation concept from HALFlow (wang2021hierarchical) into the point correlation look-up mechanism of PV-RAFT (pvraft) to learn fine-grained correlation features of sparse 4D radar points, thereby guiding the right flowing direction of point motion. Specifically, for each point $\mathbf{p}_{i,j}^{12}$ in $\mathbf{P}^{12}_{i}$, we first retrieve the initial correlation features from the correlation volume using nearest neighbor queries. Subsequently, based on the 3D local geometric structure of $\mathbf{p}_{i,j}^{12}$ and its neighboring points in $\mathbf{P}^{12}_{i}$, we adaptively weight the initial correlation features of these neighbors to obtain the final correlation feature $\mathbf{cf}_{i,j}^{12}$.
Point Motion Flow Update and Confidence Weight Prediction
@@
The motion features $\mathbf{MF}^{12}_{i}$ is used as the input to the GRU (gru) to update the current hidden state $\mathbf{h}^{12}_{i}$. Instead of predicting updates to the pose directly, we predict updates in the space of point motion flow fields. We map the updated hidden state through two additional MLP layers to produce two outputs: (1) a revision flow field $\Delta\mathbf{FL}^{12}_{i}\in\mathbb{R}^{N\times 3}$ and (2) associated confidence $\mathbf{w}^{12}_{i}\in\mathbb{R}_{+}^{N\times 3}$. The revision $\Delta\mathbf{FL}^{12}_{i}$ is a correction term predicted by the network to correct errors in the all-point correspondence field. We denote the corrected correspondence as $\mathbf{P}^{12*}_{i}=\mathbf{P}^{12}_{i}+\Delta\mathbf{FL}^{12}_{i}$. The confidence term is further processed with a sigmoid operation to constrain its values within the $(0,1)$. This term mitigates the adverse impact of inaccurate point motion flow estimation caused by occlusion, moving objects and noisy points on pose estimation accuracy.
All-point Motion-only Bundle Adjustment Layer
@@
$E(\mathbf{T}^{1},\mathbf{T}^{2})=\left\|\mathbf{P}^{12*}_{i}-\mathbf{T}^{2}\cdot(\mathbf{T}^{1})^{-1}\cdot\mathbf{P}^{1}\right\|^{2}_{\Sigma_{12}},$ (8)
where $\left\|\cdot\right\|^{2}_{\Sigma}$ is the Mahalanobis distance which weights the error terms based on the confidence weights $\mathbf{w}^{12}_{i}$. In this paper, we optimize the pose using the Gauss-Newton algorithm, based on the special Euclidean Lie algebra $\text{se}(3)=\left\{\xi=\left(\begin{array}[]{c}\rho\\ \phi\end{array}\right),\ \rho\in\mathbb{R}^{3},\ \rho\in\text{so}(3)\right\}$. First, we compute the Jacobian matrices $\mathbf{J}^{1}\in\mathbb{R}^{N\times 3\times 6}$ and $\mathbf{J}^{2}\in\mathbb{R}^{N\times 3\times 6}$ of the cost function with respect to $\mathbf{T}^{1}$ and $\mathbf{T}^{2}$ using the perturbation model.
$\mathbf{J}^{1}$ $=\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right]\circ\text{Adj}_{\mathbf{T}^{12}},$ (11)
$\mathbf{J}^{2}$ $=-\left[\begin{array}[]{cc}\mathbf{I}&-(\mathbf{R}^{12}\circ\mathbf{P}^{1}+\mathbf{t}^{12})^{\wedge}\\ \mathbf{0}^{T}&\mathbf{0}^{T}\end{array}\right],$ (14)
where $\left(\cdot\right)^{\wedge}$ denotes the antisymmetric matrix operation and $\text{Adj}_{\left(\cdot\right)}$ denotes the adjoint operator.
Then, we linearizes the cost function, and solves for the update:
$\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J}\Delta\xi=-\mathbf{J}^{T}E(\mathbf{T}^{1},\mathbf{T}^{2}),$ (15)
where $\Delta\xi=(\Delta\xi^{1},\Delta\xi^{2})$ represents the pose update in the Lie algebra. Eqn.15 can be rewritten as a linear system:
$\mathbf{H}\Delta\xi=\mathbf{b},\quad\mathbf{H}=\mathbf{J}^{T}\text{diag}(\mathbf{w}^{12}_{i})\mathbf{J},\quad\mathbf{b}=-\mathbf{J}^{T}E(\mathbf{T}^{1},\mathbf{T}^{2}).$ (16)
@@
Training and Implementation
DNOI-4DRO is implemented in PyTorch, with the LieTorch extension (teed2021tangent) used to perform backpropagation in the tangent space of all group elements.
Constructing Frame Graph